a ball made of a long string of thickness about 2mm the radius of the ball is 10cm esimate the length of the string and possible error inyour calculation.
1
Expert's answer
2019-09-23T09:20:17-0400

"length\ of\ string\times thicknes=area\ of\ ball"


"l\times0.2\ cm=4\pi r^2=4\times3.14\times10^2=1256" cm


"l=\frac{1256}{.2}=6280\ cm\\\ \ \ \ =62.8\ m"


Error in calculation will be

"\frac{1}{2}\times10^{-n}=0.05"


"n=1\ here\\\ because\ no.\ of\ digit\ after\ decimal\ is\ one"


error in calculation="0.05 \ m"